A weekly business magazine offers a 1-yearsubscription for $58 and a 3-year subscription for $116. If



you thought you would read the magazine for at least the next 3 years, and consider 20% as a minimum



rate of return, which way would you purchase the magazine: With three 1-year subscriptions or a single



3-year subscription?

1
Expert's answer
2021-11-22T08:10:03-0500

Let us assume that payments are made in the beginning of period

Present worth of 1 year subscription option=45*(P/A,0.10,6)*(1+0.10)

Let us calculate the interest factor




Present worth of 1 year subscription option=45*4.3553*1.10=$215.59

Present worth of 3 year subscription option=118+118*(P/F,0.10,3)

Let us calculate the interest factor

(P/F,0.10,3)=1/(1+0.10)^3=0.7513

Present worth of 3 year subscription option=118*+118*0.7513=$206.65


We can observe that 3-year option is cheaper. I shall prefer it.